OpenStudy (anonymous):

HELP PLEASE. What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57

OpenStudy (anonymous):

Can you solve an equation like this? 2x+1=7

OpenStudy (anonymous):

yes. x = 3

OpenStudy (anonymous):

Yes. So this is the same thing. Except now you need to isolate ONE variable.(Bring it on one side) because you can only solve equations with ONE variable.

OpenStudy (anonymous):

So: 7x = 21 + 2y x = (21+2y)/7

OpenStudy (anonymous):

i thought you subtract 7x - 4x

OpenStudy (anonymous):

now you have X by itself. So then, put the NEW x equation (the one we just made) into the OTHER equation. So: 4x + y = 57 4((21+2y)/7) + y = 57

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

why subtract ?

OpenStudy (anonymous):

i have to do one equation at a time, i didnt know that now i understand i do the first equation then put it in to the second

OpenStudy (anonymous):

i got 15 as the answer

OpenStudy (anonymous):

Let me checl

OpenStudy (anonymous):

The answer should be different. Try it again

OpenStudy (anonymous):

When you find the right values for x and y, you will find that the equations are both true.